OpenStudy (anonymous):

Convert 150 g/L to the unit g/mL.

OpenStudy (anonymous):

\[\frac{ 150g }{ 1 } * \frac{ 1 }{ 1000 } = \frac{ 150 }{ 1000 } = 0.150g/mL\]

OpenStudy (anonymous):

150g/L = 150g/1000 mL = 150x10^-3 = 15x10^-2 = 0,0

OpenStudy (anonymous):

= 0.15 :) sorry